# Contact Models in Robotics: a Comparative Analysis

### v2

@@ II Rigid contact modelling
where $\Phi:\mathbb{R}^{n_{q}}\mapsto\mathbb{R}^{m}$ is an holonomic constraint function of dimension $m$, which depends on the nature of the constraint. For solving, it is more practical to proceed to an index reduction [53] by differentiating (3) w.r.t. time, in order to express it as a constraint on joint velocities:
@@
In the case of rigid bodies — purely inelastic impacts, and exact collision detection, i.e. $\Phi(q)_{N}=0$ — the reference velocity $c_{N}^{*}$ can be set to 0. However, setting this velocity to a non-null quantity may be useful to improve modeling on both physical and numerical aspects. A first benefit is the possibility of accounting for impacts that may occur when two objects collide with non-null relative normal velocity. The most common impact law stipulates $c^{*}=-ec^{t}$ where $e$ is the restitution coefficient, which adjusts the quantity of energy dissipated during the collision. When time-stepping methods are employed, one cannot avoid penetration errors, i.e. $\Phi(q)_{N}<0$, without using stabilization by reprojection techniques [65] which are computationally expensive to use in robotics due to the cost of detecting a collision. However, it is still possible to prevent these errors from dramatically growing over time via a Baumgarte correction [66] which sets $c^{*}=k_{B}\max(0,-\Phi(q^{t})_{N})$ where the Baumgarte coefficient $k_{B}$ is set to be proportional to $\frac{1}{\Delta t}$.
In addition, in many cases in robotics, Delassus’ matrix $G$ is rank deficient. Such physical systems are said to be hyperstatic, and because $\mathrm{rank}(J)>n_{v}$, several $\lambda$ values may lead to the same trajectory. This under-determination can be circumvented by relaxing the rigid-body hypothesis, e.g. the Signorini condition, and considering compliant contacts via a reference velocity linearly depending on $\lambda$ as represented in Fig. 3. Indeed, with $c^{*}=-R\lambda$ where $R$ is a diagonal matrix with non-null and positive elements only on the normal components, called compliance and whose value is a property of the material, the original Delassus matrix $G$ is replaced by the damped matrix $\tilde{G}=G+R$ which is full rank. At this stage, one should note that the physical compliance acts on the conditioning in an equivalent way to a numerical regularization.
@@
Combining the Coulomb’s law for friction with the Signorini condition evoked earlier, we finally get three distinct cases corresponding to a sticking contact point (13a), a sliding contact point (13c) or a take-off (13b):
@@
At this point, it is worth mentioning that the problem (15), which we refer to as NCP, does not derive from a convex optimization problem, thus making its solving complex. Alternatively, one can see the frictional contact problem as two interleaved convex optimization problems [69, 70, 36, 71, 6] whose unknowns, $\lambda$ and $v$, appear in both. Other formulations exist and we refer to [68] for a more complete review on the NCP. Practically, the non-convexity can induce the existence of multiple, or even an infinite number of contact forces satisfying (15). As mentioned earlier, this can be due to normal forces, but tangential components can also cause under-determination (Fig. 4). In this situation, it would be preferable for a simulator to provide the minimum norm solution in forces. This property can prevent a simulator from exhibiting internal friction forces compressing or stretching the objects (Fig. 4, right). Indeed, such forces would not coincide with the forces observed by force sensors and would rather correspond to some internal deformations of the objects, which should thus be considered soft and no more rigid. In the following, we will use the term “internal forces” to denote the force component deviating from the minimum norm solution. Additionally, these internal forces might also be problematic as it is impossible to characterize them. This may induce inconsistent derivatives, which become critical in the context of differentiable simulation.
@@
More precisely, at each time step, a simulator must first detect which geometries are colliding and compute their separation vector $\Phi$. The GJK and EPA algorithms are widely adopted for their low computational cost. HPP-FCL [45, 46, 47], an extension of the Flexible Collision Library (FCL) [72] and libccd [73] implement them efficiently. Some simulators such as Bullet [30] or ODE [29] also re-implement the same algorithm as an internal routine.
@@
Figure 5: Simulation routines. When simulating rigid bodies with frictional contacts, a physics engine goes through a sequence of potentially challenging sub-problems: collision detection, collision resolution, and integration time step.
@@
Collision resolution
@@ III-A Linear Complementarity Problem
Projected Gauss-Seidel. Due to its easy implementation and its early stopping property, the projected Gauss-Seidel (PGS) algorithm (Alg. 1) algorithm represents an attractive alternative and was widely adopted as the default solver by many physics engines, such as in Bullet [30], PhysX [75], ODE [29], and DART [31, 7] simulators. This iterative algorithm loops on contact points and successively updates the normal and tangent contact forces. Because PGS works separately on each contact point, the update compensates for the current errors due to the estimated forces from other contact points. Yet, as illustrated in the experimental section IV, this process induces the emergence of internal forces during the solving. Moreover, Gauss-Seidel-based approaches are similar to what is also known as block coordinate descent in the optimization literature. As first-order algorithms, they do not benefit from improved convergence rates or robustness with respect to their conditioning, unlike second-order algorithms. In parallel, the linearization of the second-order cone causes the loss of the isotropy for friction, as stated by Coulomb’s law. By choosing the axes for the facets of the pyramid and due to the maximum dissipation principle, it is established that one incidentally biases the friction forces towards the corners [84, 85], as illustrated in Fig. 6. This error is sometimes mitigated by increasing the number of facets, which also comes at the cost of more computations.
@@ 6             $\lambda^{(i)}_{T}\leftarrow\text{clamp}(\lambda^{(i)}_{T},\mu_{i}\lambda_{N})$;
PGS. The PGS algorithm can be directly adapted to handle the CCP problem [76] (Alg. 2). In the light of what follows, the algorithm even becomes equivalent to a projected gradient descent which is a classical constrained optimization technique.
@@ 4             $\lambda^{(i)}\leftarrow\text{proj}_{K_{\mu_{i}}}(\lambda^{(i)})$;
Once the contact problem is formulated as an optimization problem, any optimization algorithms can be employed to solve it and classical optimization theory provides convergence guarantees. Here, we propose to study an ADMM algorithm [87] to solve (20), but Interior Point algorithms [88] could also be used [32]. A benefit of using the family of proximal algorithms like ADMM is their natural ability to handle the numerical issues coming from ill-conditioned and hyper-static cases [89, 90]. This property makes it possible to accurately simulate hard contacts, i.e., without any shift due to compliance $R=0$, and is reported in Tab. III by the "hard contacts" column. Another by-product of such methods is the implicit regularization they induce on the found solution, which removes the potential internal forces. This last property is an empirical observation resulting from the experimental section IV and, to our knowledge, has not yet been proven by the literature of proximal optimization. Therefore, it remains to be confirmed by subsequent work.
@@ 5       $\gamma\leftarrow\gamma+\rho(\lambda-z)$;
Both Drake and MuJoCo use a Newton solver to tackle (22) (Alg. 4). Due to the non-linearity of the second term of (22), this approach requires updating the inverse of the Hessian at every iteration (Alg. 4,line 4). As proposed in [35], the use of advanced algebra routines allows to reduce the computational burden of each step. In this work, we provide an implementation of the Newton algorithm with an Armijo backtracking linesearch (using parameters from [35]), which is close to what is done in the MuJoCo and Drake simulators.
@@ III-D Tackling the NCP
PGS. The PGS algorithm exploited for LCP and CCP problems can easily be adapted to the NCP case by changing the clamping step (Alg. 1,line 1) or the normal projection (Alg. 2, line 2) for a horizontal projection on the cone (Alg. 6, line 6). However, it is worth noting that such approaches have fewer convergence guarantees than their relaxed counterpart [27]. As with every Gauss-Seidel approach, the methods inherited from the sensitivity to ill-conditioning and spurious internal forces.
@@ III-E Implementation details
Warm-starting the solver by providing the contact forces from the previous time step allows to greatly reduce the required computation. Indeed, in the case of a persisting contact between two objects, the contact forces are being cached and reused as an initial guess when solving for the contact forces of the next time step. This relies on the ability of the contact solver algorithm to be warm-started. This excludes Interior Point [88] algorithms, as they would only benefit from an initial guess close to the so-called central path [89]. By contrast, the feasible set of contact forces may change from one time step to the other, even in the case of a persisting contact point. On the opposite, ADMM and, more generally, Augmented Lagrangian (AL) methods can naturally be warm started: not only the primal (i.e., contact forces) and dual (i.e., contact velocities) variables, but also the proximal parameter is initialized with the previous values.
@@
Proximal parameter adaptation. In the context of ADMM (Alg. 3), the algorithm from [54] can also be favorably combined with the adaptation of the proximal parameter. Indeed, updating the regularized Cholesky can be done at almost no cost by using [54]. In our implementation, we follow the work from [93] to detect when and how $\rho$ should be adapted. More precisely, whenever the primal residual is significantly greater than the dual one (a threshold for the ratio has to be set, a typical value being 10), $\rho$ should be increased in order to better enforce the constraint of the problem and thus, reduce the primal residual. The proximal parameter $\rho$ is then updated via a spectral rule which multiplies it by $\kappa^{0.05}$, $\kappa$ being the condition number of the Delassus matrix, defined as the ratio between the largest and the smallest eigenvalues. Conversely, whenever the dual residual dominates the primal one, $\rho$ should be decreased by dividing it with the same factor. Alternatively, we could use a linear update rule for $\rho$ as it is done in OSQP [94] which would be less efficient in the case of ill-conditioned problems.
@@ IV Experiments
In this section, we evaluate the performances and behaviors of the formulations explained in Sec. III. To fairly compare and benchmark the various algorithmic formulations, we have implemented them in a unified C++ framework called ContactBench. In the following, we denote by RaiSim and Drake our re-implementation of the contact solvers described in the corresponding papers [33, 35]. For Drake, it is worth noting that our implementation uses a backtracking line-search with an Armijo condition instead of the line-search proposed in the original paper [35]. Our framework extensively relies on the Pinocchio library [24] for rigid body algorithms and HPP-FCL [45, 56] implementation of GJK and EPA for collision detection. ContactBench will be released as open-source upon article acceptance.
@@
Except where expressly indicated, we use the following values for the solvers parameters: $\epsilon_{abs}=10^{-6}$, $n_{iter}=10^{4}$ and $\Delta t=1ms$. For the ADMM algorithm, the proximal parameter $\rho$ is adapted dynamically as previously detailed (Sec. III-E).
@@ IV-A Evaluation of physical correctness
Underdetermined contact problems. Underdetermination occurs when infinite combinations of contact forces lead to the same trajectory. These artifacts happen on the normal and tangential components of contact forces, as depicted in Fig. 11. As shown in Fig. 11, the solution found depends on the numerical scheme. We observe that the per-contact approaches (Alg. 2,5 and 6) exhibit spurious internal forces at stiction, values which are not controlled by the algorithms. On the opposite, the algorithms working directly on the global contact problem with a proximal regularization (Alg. 3 and 7) seem to avoid injecting such artifacts in the contact forces (Fig. ,11). As future work, it would be interesting to investigate the theory behind the latter conjecture.
Figure 11: Applying a linearly growing force along the x-axis to a cube on a plane. Left: The cube is at stiction before it starts sliding after approximately 0.25s. The tangential velocity differs depending on the contact model: CCP induces some bounces which occasionally dissipate energy during impacts, while RaiSim violates the MDP leading to contact points sliding faster than in the case of NCP. Right: At stiction, multiple combinations of tangential forces may lead to the same trajectory. There are four curves for each contact model, each curve accounting for the y-component from one of the four contact forces on the cube. Gauss-Seidel-like solvers, e.g. RaiSim and PGS, exhibit internal forces "stretching" the cube at stiction before the MDP enforces these forces to disappear when the cube starts to slide. RaiSim relaxes the maximum dissipation principle so the friction forces are not opposed to the movements, and internal forces persist when the cube is sliding. Eventually, ADMM avoids injecting spurious internal forces even at stiction.
@@
Robustness to ill-conditioned contact problems. More generally, the contact problem becomes challenging when the ratio between the biggest and the smallest eigenvalue of the Delassus matrix grows. The experiment of Fig. 12 exhibits the convergence issues of per-contact approaches when simulating systems with a strong coupling between the different contact points, which causes large off-diagonal terms on the matrix $G$. In this situation, the latter approaches hit the maximum number of iteration before convergence, leading to unrealistic trajectories. Such a behavior can be expected as supposing the matrix to be diagonally dominant is a classical hypothesis ensuring the convergence of Gauss-Seidel methods. On the contrary, the proximal algorithms account for off-diagonal terms of $G$, and only rely on a regularized inverse of $G$ (Alg. 3, line 3), and thus robustly converge towards an optimal solution.
Figure 12: Simulation of ill-conditioned systems. Left: Stacking a heavy cube ($10^{3}$kg) on a light one ($10^{-3}$kg) makes the problem ill-conditioned and, therefore, not solvable via per-contact algorithms (CCP/PGS, NCP/PGS and RaiSim) which results in the violation of the contact complementarity criterion (15). By contrast, the ADMM, staggered projections and Newton approaches appear to be robust in this case. Right: The accuracy of the simulators improves when the ratio between the masses of the two cubes gets close to one. The ADMM and Newton algorithms are less affected by this ratio than PGS.
@@
Figure 13: Simulation of Solo-12 with varying compliance for the contacts with the floor. Left: Adding a compliance to the contact model relaxes the Signorini condition. Center: This compliance also relaxes the Coulomb’s law of friction. Right: Compliance also regularizes the problem, removing the spurious internal forces in the under-determined cases.

### v3

@@ II Rigid contact modelling
where $\Phi:\mathbb{R}^{n_{q}}\mapsto\mathbb{R}^{m}$ is a holonomic constraint function of dimension $m$, which depends on the nature of the constraint. For solving, it is more practical to proceed to an index reduction [53] by differentiating (3) w.r.t. time, in order to express it as a constraint on joint velocities:
@@
In the case of rigid bodies, the reference velocity $c_{N}^{*}$ can be set to $\frac{\Phi(q^{t})}{\Delta t}$ to complete the Taylor expansion of (6). However, adding bias terms to this velocity may be useful to improve modeling on both physical and numerical aspects. A first benefit is the possibility of accounting for impacts that may occur when two objects collide with non-null relative normal velocity. The most common impact law stipulates to introduce a bias term $-ec^{t}$ where $e$ is the restitution coefficient, which adjusts the quantity of energy dissipated during the collision. When time-stepping methods are employed, one cannot avoid penetration errors, i.e. $\Phi(q)_{N}<0$, without using stabilization by reprojection techniques [65] which are computationally expensive to use in robotics due to the cost of detecting a collision. However, it is still possible to prevent these errors from dramatically growing over time via a Baumgarte correction [66] which adds $k_{B}\max(0,-\Phi(q^{t})_{N})$ to the reference velocity $c^{*}$ and where the Baumgarte coefficient $k_{B}$ is set to be proportional to $\frac{1}{\Delta t}$.
In addition, in many cases in robotics, Delassus’ matrix $G$ is rank deficient. Such physical systems are said to be hyperstatic, and because $\mathrm{rank}(J)>n_{v}$, several $\lambda$ values may lead to the same trajectory. This under-determination can be circumvented by relaxing the rigid-body hypothesis, e.g. the Signorini condition, and considering compliant contacts via a reference velocity linearly depending on $\lambda$ as represented in Fig. 3. Indeed, by adding $-R\lambda$ to $c^{*}$ where $R$ is a diagonal matrix with non-null and positive elements only on the normal components, called compliance and whose value is a property of the material, the original Delassus matrix $G$ is replaced by the damped matrix $\tilde{G}=G+R$ which is full rank. At this stage, one should note that the physical compliance acts on the conditioning in an equivalent way to a numerical regularization.
@@
The Non-linear Complementarity Problem. Combining the Coulomb’s law for friction with the Signorini condition evoked earlier, we finally get three distinct cases corresponding to a sticking contact point (13a), a sliding contact point (13c) or a take-off (13b):
@@
At this point, it is worth mentioning that the problem (15), which we refer to as NCP, does not derive from a convex optimization problem, thus making its solving complex. Alternatively, one can see the frictional contact problem as two interleaved convex optimization problems [69, 70, 36, 71, 6] whose unknowns, $\lambda$ and $v$, appear in both. Other formulations exist and we refer to [68] for a more complete review on the NCP. Practically, the non-convexity can induce the existence of multiple, or even an infinite number of contact forces satisfying (15). As mentioned earlier, this can be due to normal forces, but tangential components can also cause under-determination (Fig. 4). In this situation, it would be preferable for a simulator to provide the minimum norm solution in forces. This property can prevent a simulator from exhibiting internal friction forces compressing or stretching the objects (Fig. 4, right). Indeed, such forces may not coincide with the forces observed by force sensors and would rather correspond to some internal deformations of the objects, which should thus be considered soft and no more rigid. In the following, we will use the term “internal forces” to denote the force component deviating from the minimum norm solution. Additionally, these internal forces might also be problematic as it is difficult to characterize them. This may induce inconsistent derivatives, which become critical in the context of differentiable simulation.
@@
More precisely, at each time step, a simulator must first detect which geometries are colliding and compute their separation vector $\Phi$. The GJK and EPA algorithms are widely adopted for their low computational cost. HPP-FCL [45, 46, 47], an extension of the Flexible Collision Library (FCL) [72] and libccd [73] implement them efficiently. Some simulators such as Bullet [30], ODE [29] or PhysX [74] also re-implement the same algorithm as an internal routine.
@@
Figure 5: Simulation routines. When simulating rigid bodies with frictional contacts, a physics engine goes through a sequence of potentially challenging sub-problems: collision detection, contact forces computation, and integration time step.
@@
Drake [34] BSD3 C++, Python
Forces computation
@@ III-A Linear Complementarity Problem
Projected Gauss-Seidel. Due to its easy implementation and the possibility to early-stop it, the projected Gauss-Seidel (PGS) algorithm (Alg. 1) algorithm represents an attractive alternative and was widely adopted as the default solver by many physics engines, such as in Bullet [30], PhysX [74], ODE [29], and DART [31, 7] simulators. This iterative algorithm loops on contact points and successively updates the normal and tangent contact forces. Because PGS works separately on each contact point, the update compensates for the current errors due to the estimated forces from other contact points. Yet, as illustrated in the experimental section IV, this process induces the emergence of internal forces during the solving. Moreover, Gauss-Seidel-based approaches are similar to what is also known as block coordinate descent in the optimization literature. As first-order algorithms, they do not benefit from improved convergence rates or robustness with respect to their conditioning, unlike second-order algorithms.
In parallel, the linearization of the second-order cone causes the loss of the isotropy for friction, as stated by Coulomb’s law. By choosing the axes for the facets of the pyramid and due to the maximum dissipation principle, it is established that one incidentally biases the friction forces towards the corners [84, 85], as illustrated in Fig. 6. This error is sometimes mitigated by increasing the number of facets, which also comes at the cost of more computations.
@@ 6             $\lambda^{(i)}_{T}\leftarrow\text{clamp}(\lambda^{(i)}_{T},\mu_{i}\lambda_{N})$;
PGS. The PGS algorithm can be directly adapted to handle the CCP problem [76] (Alg. 2) but inherits from the first-order convergence rates ([76] exhibits in the order of hundreds of iterations to converge in general). In the light of what follows, the algorithm even becomes equivalent to a projected gradient descent which is a classical constrained optimization technique.
@@ 4             $\lambda^{(i)}\leftarrow\text{proj}_{K_{\mu_{i}}}(\lambda^{(i)})$;
Once the contact problem is formulated as an optimization problem, any optimization algorithms can be employed to solve it and classical optimization theory provides convergence guarantees. Here, we propose to study an ADMM algorithm [87], an advanced first-order algorithm known to be efficient to reach mild accuracy and which can stall when further improving the solution. As pointed out in [32], Interior Point algorithms [88] could also be used to reach higher precision solutions even though we do not find them in any of the robotics simulators here mentioned. A benefit of using the family of proximal algorithms like ADMM is their natural ability to handle the numerical issues coming from ill-conditioned and hyper-static cases [89, 90]. This property makes it possible to accurately simulate hard contacts, i.e., without any shift due to compliance $R=0$, and is reported in Tab. III by the "hard contacts" column. Another by-product of such methods is the implicit regularization they induce on the found solution, which removes the potential internal forces. This last property is an empirical observation resulting from the experimental section IV and, to our knowledge, has not yet been proven by the literature of proximal optimization. Therefore, it remains to be confirmed by subsequent work.
@@ 5       $\gamma\leftarrow\gamma+\rho(\lambda-z)$;
Both Drake and MuJoCo use a Newton solver to tackle (22) (Alg. 4). Due to the non-linearity of the second term of (22), this approach requires updating the inverse of the Hessian at every iteration (Alg. 4,line 4). As proposed in [35], the use of advanced algebra routines allows to reduce the computational burden of each step. In this work, we provide an implementation of the Newton algorithm with an Armijo backtracking line search (using parameters from [35]). MuJoCo and Drake additionally implement an exact line search which improves performance and this difference should be kept in mind when interpreting the results obtained with our implementation.
@@ III-D Tackling the NCP
PGS. The PGS algorithm exploited for LCP and CCP problems can easily be adapted to the NCP case by changing the clamping step (Alg. 1,line 1) or the normal projection (Alg. 2, line 2) for a horizontal projection on the cone (Alg. 6, line 6). However, it is worth noting that such approaches have fewer convergence guarantees than their relaxed counterpart [27]. As with every Gauss-Seidel approach, the methods inherited from the sensitivity to ill-conditioning and jamming internal forces.
@@ III-E Implementation details
Warm-starting the solver by providing the contact forces from the previous time step allows to greatly reduce the required computation. Indeed, in the case of a persisting contact between two objects, the contact forces are being cached and reused as an initial guess when solving for the contact forces of the next time step. This relies on the ability of the contact solver algorithm to be warm-started. This excludes Interior Point [88] algorithms, as they would only benefit from an initial guess close to the so-called central path [89]. By contrast, the feasible set of contact forces may change from one time step to the other, even in the case of a persisting contact point. On the opposite, ADMM and, more generally, Augmented Lagrangian (AL) methods can naturally be warm started: not only the primal (i.e., contact velocities) and dual (i.e., contact forces) variables, but also the proximal parameter is initialized with the previous values.
@@
Proximal parameter adaptation. In the context of ADMM (Alg. 3), the algorithm from [54] can also be favorably combined with the adaptation of the proximal parameter. Indeed, updating the regularized Cholesky can be done at almost no cost by using [54]. In our implementation, we follow the work from [93] to detect when and how $\rho$ should be adapted. More precisely, whenever the primal residual is significantly greater than the dual one (a threshold for the ratio has to be set, a typical value being 10), $\rho$ should be increased in order to better enforce the constraint of the problem and thus, reduce the primal residual. The proximal parameter $\rho$ is then updated via a spectral rule which multiplies it by $\kappa^{0.05}$, $\kappa$ being the condition number of the Delassus matrix, defined as the ratio between the largest and the smallest eigenvalues. Conversely, whenever the dual residual dominates the primal one, $\rho$ should be decreased by dividing it with the same factor. The condition number $\kappa$ can be efficiently evaluated beforehand via a power-iteration algorithm whose iterates have a computational cost equivalent to the ones from the ADMM algorithm. This procedure is detailed and evaluated in [93]. Alternatively, we could use a linear update rule for $\rho$ as it is done in OSQP [94] which would be less efficient in the case of ill-conditioned problems.
@@ IV Experiments
In this section, we evaluate the performances and behaviors of the formulations explained in Sec. III. To fairly compare and benchmark the various algorithmic formulations, we have implemented them in a unified C++ framework called ContactBench. In the following, we denote by RaiSim and Drake our re-implementation of the contact solvers described in the corresponding papers [33, 35]. For Drake, it is worth noting that our implementation uses a backtracking line-search with an Armijo condition instead of the line-search proposed in the original paper [35]. Our framework extensively relies on the Pinocchio library [24] for rigid body algorithms and HPP-FCL [45, 56] implementation of GJK and EPA for collision detection. Our code is made open-source in the Contactbench C++ library (https://github.com/Simple-Robotics/contactbench).
@@
Except where expressly indicated, we use the following values for the solvers’ parameters: an absolute convergence criterion $\epsilon_{abs}=10^{-6}$, a maximum number of iteration $n_{iter}=10^{4}$ and a time-step $\Delta t=1ms$. For the ADMM algorithm, the proximal parameter $\rho$ is adapted dynamically as previously detailed (Sec. III-E).
@@ IV-A Evaluation of physical correctness
Figure 11: Applying a linearly growing force along the x-axis to a cube on a plane. The cube has a mass of 1kg, a side length of 0.2m, a friction coefficient of 0.4 and the external force grows linearly from 0 to 20N over 1s. This induces relatively high velocities in a robotic context that are useful for illustrative purposes here.
@@
In Figure 11, we simulate a cube dragged on a plane and measure the integral error between the trajectory obtained with the CCP/ADMM solver with various $\Delta t$ and a reference one computed via the NCP/PGS solver with small time step $\Delta t=10^{-5}s$. The deviation from the reference trajectory is quantified via the integral consistency error defined as $\sum_{\tau=0}^{T}\|q^{\tau}-\overline{q}^{\tau}\|\Delta t$. As detailed in Sec. III-B, the error between the trajectories obtained with the CCP and NCP models is proportional to the time step $\Delta t$.
Underdetermined contact problems. Underdetermination occurs when infinite combinations of contact forces lead to the same trajectory. These artifacts happen on the normal and tangential components of contact forces, as depicted in Fig. 12. As shown in Fig. 12, the solution found depends on the numerical scheme. We observe that the per-contact approaches (Alg. 2,5 and 6) exhibit jamming internal forces at stiction, values which are not controlled by the algorithms. On the opposite, the algorithms working directly on the global contact problem with a proximal regularization (Alg. 3 and 7) seem to avoid injecting such artifacts in the contact forces (Fig. ,12). As future work, it would be interesting to investigate the theory behind the latter conjecture.
Figure 12: A cube is dragged on a plane along the x-axis similarly to the case studied in Fig. 11. Left: The cube is at stiction before it starts sliding after approximately 0.25s. The tangential velocity differs depending on the contact model e.g. RaiSim violates the MDP leading to contact points sliding faster than in the case of NCP. Right: At stiction, multiple combinations of tangential forces may lead to the same trajectory. There are four curves for each contact model, each curve accounting for the y-component from one of the four contact forces on the cube. Gauss-Seidel-like solvers, e.g. RaiSim and PGS, exhibit internal forces "stretching" the cube at stiction before the MDP enforces these forces to disappear when the cube starts to slide. RaiSim relaxes the maximum dissipation principle so the friction forces are not opposed to the movements, and internal forces persist when the cube is sliding. Eventually, ADMM avoids injecting jamming internal forces even at stiction.
@@
Robustness to ill-conditioned contact problems. More generally, the contact problem becomes challenging when the ratio between the biggest and the smallest eigenvalue of the Delassus matrix grows. The experiment of Fig. 13 exhibits the convergence issues of per-contact approaches when simulating systems with a strong coupling between the different contact points, which causes large off-diagonal terms on the matrix $G$. In this situation, the latter approaches hit the maximum number of iterations before convergence, leading to unrealistic trajectories. Such a behavior can be expected as supposing the matrix to be diagonally dominant is a classical hypothesis ensuring the convergence of Gauss-Seidel methods. On the contrary, the proximal algorithms account for off-diagonal terms of $G$, and only rely on a regularized inverse of $G$ (Alg. 3, line 3), and thus robustly converge towards an optimal solution.
Figure 13: Simulation of ill-conditioned systems. Left: Stacking a heavy cube ($10^{3}$kg) on a light one ($10^{-3}$kg) makes the problem ill-conditioned and, therefore, not solvable via per-contact algorithms (CCP/PGS, NCP/PGS and RaiSim) which results in the violation of the contact complementarity criterion (15). By contrast, the ADMM, staggered projections, and Newton approaches appear to be robust in this case. Right: The accuracy of the simulators improves when the ratio between the masses of the two cubes gets close to one. The ADMM and Newton algorithms are less affected by this ratio than PGS.
@@
Figure 14: Simulation of Solo-12 with varying compliance for the contacts with the floor. The robot is in a standing position and perturbed with an external horizontal force. Left: Adding a compliance to the contact model relaxes the Signorini condition. Center: This compliance also relaxes the Coulomb’s law of friction. Right: Compliance also regularizes the problem, removing the jamming internal forces in the under-determined cases.